OpenStudy (anonymous):

What is equal to (9b^2-b+4)/(3b-2) - (2b^2+b-8)/(3b-2)?

OpenStudy (anonymous):

are there any choices?

OpenStudy (anonymous):

7b^2+12/3b-2 7b^2-2b+12/3b-2 7b^2-4/3b-2 11b^2-2b+12/3b-2

OpenStudy (anonymous):

You'll notice both terms have a common denominator, so all you need to do is subtract like terms.\[\frac{9b^2-2b^2-b-b+4+8}{3b-2}\]

OpenStudy (anonymous):

All I did was move some stuff around; I'll leave the computation to you.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

My pleasure.

OpenStudy (anonymous):

Would that make the answer 7b^2+12/3b-2?

OpenStudy (anonymous):

you're missing a term in your numerator; what happened to the b's?

OpenStudy (anonymous):

\[-b-b\neq 0\]

OpenStudy (anonymous):

I thought they would cancel out so that means that its 7b^2-2b+12/3b-2?

OpenStudy (anonymous):

yeah be careful that you check the sign of a term when you're adding them\[+b+b=2b\]\[+b-b=0\]\[-b-b=-2b\]

OpenStudy (anonymous):

Okay, that makes sense. Thank you!

OpenStudy (anonymous):

No problem.